OpenStudy (anonymous):

Find the time, t, in seconds it takes for an object to fall from the top of a 1600 ft building by using the function (hint: set h=0).

OpenStudy (anonymous):

the function is h=-16t^2+1600

OpenStudy (anonymous):

h is height??

OpenStudy (amistre64):

\[mgh=\frac12mv^2\] \[v=\sqrt{2gh}\] \[\sqrt{2gh}=gt\] \[\frac{\sqrt{2gh}}{g}=t\] maybe

OpenStudy (amistre64):

or simply; when t^2 = 100 :/